### 22 Feb 2024 v1

@@ 1 Introduction
  • We further propose IR-DRO, a principled optimization-based algorithm that utilizes the instance reweighting formulation to dynamically identify and prioritize crucial and informative samples for the continual training of LLMs.
  • Extensive experiments conducted on various pre-trained models, such as OPT, Sheared-LLaMA, and LLaMA, have confirmed the effectiveness of our proposed algorithms. Our methods achieve improved performance on evaluation benchmarks after continual pre-training and instruction tuning, respectively.
@@ 2.1 Pretraining of LLMs
2.2 Data Re-weighting and Selection in NLPs
It has been known that i.i.d sampling may not be the best strategy when training with large and noisy data potentially coming from compositional domains. DoReMi (Xie et al., 2023) proposes using an auxiliary model to determine the optimal weights for different domain data and achieve better performance. Skill-It (Chen et al., 2023) introduces an online data sampling technique that dynamically updates the mixture of data corpus from predefined skill sets. Some seek methods with a finer granularity, i.e., reweighting data at the instance level. These works generally use the loss values (Loshchilov & Hutter, 2015; Jiang et al., 2019) or the gradient norms (Katharopoulos & Fleuret, 2018) to derive a probability distribution for data sampling. However, these methods often require a second forward-backward process that damages the scalability of the methods (Fan & Jaggi, 2023). Fan & Jaggi (2023) uses an auxiliary model and defines a learnability score for sample reweighting, yet the auxiliary model needs to undergo a similar pretraining process, which incurs additional training costs.
@@ 4.2 Formulation
We propose a novel approach, named Instance-Reweighted Distributionally Robust Optimization (IR-DRO), that overcomes the limitations of existing methods while offering several key advantages:
@@
Specifically, we formulate a minimax optimization problem with a KL-divergence regularizer as follows:

### 01 Mar 2024 v2

@@ 1 Introduction
  • We further adapt and develop IR-DRO, a principled optimization-based algorithm that utilizes the instance reweighting formulation to dynamically identify and prioritize crucial and informative samples for the continual training of LLMs.
  • Extensive experiments conducted on various pre-trained models, such as OPT (Zhang et al., 2022), Sheared-LLaMA (Xia et al., 2023), and LLaMA (Touvron et al., 2023), have confirmed the effectiveness of our proposed algorithms. Our methods achieve improved performance on evaluation benchmarks after continual pre-training and instruction tuning, respectively.
@@ 2.1 Pretraining of LLMs
2.2 Data Re-weighting and Selection
It has been known that i.i.d sampling may not be the best strategy when training with large and noisy data potentially coming from compositional domains. DoReMi (Xie et al., 2023) proposes using an auxiliary model to determine the optimal weights for different domain data and achieve better performance. Skill-It (Chen et al., 2023) introduces an online data sampling technique that dynamically updates the mixture of data corpus from predefined skill sets. Some seek methods with a finer granularity, i.e., reweighting data at the instance level. These works generally use the loss values (Loshchilov & Hutter, 2015; Jiang et al., 2019) or the gradient norms (Katharopoulos & Fleuret, 2018) to derive a probability distribution for data sampling. However, these methods often require a second forward-backward process that damages the scalability of the methods. Fan & Jaggi (2023) uses an auxiliary model and defines a learnability score for sample reweighting, yet the auxiliary model needs to undergo a similar pretraining process, which incurs additional training costs. The assignment of weights to individual samples can also be conceptualized within the context of optimization problems. A special family of these problems, the KL-regularized distributionally robust optimization (DRO) in a min-max form, can be efficiently addressed by transforming into simpler compositional minimization forms (Qi et al., 2021; 2022). Follow-up works have demonstrated that such a formulation can mitigate issues related to data imbalance (Qi et al., 2023) and enhance contrastive learning by customizing temperatures for individual samples (Qiu et al., 2023).
@@ 4.2 Formulation
We adapt an optimization framework, named Instance-Reweighted Distributionally Robust Optimization (IR-DRO), based on the formulation of distributionally robust optimization (Qi et al., 2021; 2022). It overcomes the limitations of existing methods while offering several key advantages:
@@
Specifically, we formulate a minimax optimization problem with a KL-divergence regularizer inspired by Qi et al. (2021) as follows: